A well 28.8m deep and of diameter 2m is dug up. The soil dug up is spread all around the well to make a platform 1m high considering the fact that the loose soil settles to a height in the ratio 6:5. Find the width of the platform.

Grade:10

1 Answers

Arun
6 years ago
Dear Aryan
 
Volume of soil dug out = \pi (1)² (28.8) m³
Volume of soil that should be settled on the platform is 
5\pi * 28.8/6 m³
 
Let the width of platform be r cm.
\pi(1+r)² - (1)² (1) = 5\pi*28.8/6
(1+r)² -1 = 24
(1+r)² = 25
1+r = 5
Hence
r = 4 cm
 
Regards
Arun (askIITians forum expert)